OpenStudy (janu16):

Find an equation of a sine function with an amplitude of 2, a period of pi and a y-intercept of -4.

OpenStudy (janu16):

Can i get help please?

OpenStudy (janu16):

@phi

OpenStudy (phi):

\[ y = A \sin\left(\frac{2 \pi}{T} x + B\right) \] A is the amplitude T is the period we still have to figure out B

OpenStudy (janu16):

This is what I have so far Y=4 sin 2(x)

OpenStudy (phi):

oh, I was thinking we needed to include a phase shift B but it looks like we need an offset, so let's use \[ y = 2 \sin 2x+ B\]

OpenStudy (janu16):

Yes

OpenStudy (janu16):

I need help finding b

OpenStudy (phi):

amplitude is 2 (they tell us this, so we have to use 2, not 4) to find the offset, we use this info: y-intercept of -4. that means the point (0, -4) use x=0 and y= -4 and solve for the offset.

OpenStudy (janu16):

Nvm the amolitude is 2

OpenStudy (phi):

yes, y = 2 sin(2x) + B to find B, set x=0 and y= -4 -4 = 2 sin(2*0)+B -4 = 2*0+B -4 = B

OpenStudy (janu16):

So y=2 sin pi(x)?

OpenStudy (phi):

that has the correct amplitude, but not the correct period nor the correct offset.

OpenStudy (janu16):

So how do you fidn the period

OpenStudy (phi):

\[ y = A \sin\left(\frac{2 \pi}{T} x \right)+B\] if you know T, plug it in and simplify if you have , for example sin(x) which is sin(1*x) then 2pi/T = 1 (the number in front of the x) and solve for T to get T= 2pi. in other words, sin(x) has a period of 2pi

OpenStudy (janu16):

So y=2sin2pi(x)-4?

OpenStudy (phi):

almost. they say T=2 plug in T=2 into 2pi/T to get the number in front of the x

OpenStudy (phi):

*T = pi (not 2)

OpenStudy (janu16):

2?

OpenStudy (phi):

yes

OpenStudy (janu16):

So it would be y= 2 sin 2(x)-4?

OpenStudy (janu16):

The equation

OpenStudy (phi):

I would write it as y= 2 sin(2x) - 4

OpenStudy (janu16):

Ok thanks!

OpenStudy (janu16):

So if i had a similar problem but they gave me a phase shift instead if y intercept, how would I find B?

OpenStudy (janu16):

@phi